It is a kite. The perimeter is 6m. Each long side is 0.5m longer than each short side. How long is each side?

Answer by macston(5194)

S=short side; L=long side=S+0.5m; P=perimeter=2(L+S)
P=2(L+S) Substitute for L.
6m=2(S+0.5m+S) Divide each side by 2.
3m=2S+0.5m Subtract 0.5m from each side.
2.5m=2S Divide each side by 2.
1.25m=S ANSWER1: The short sides are 1.25 meters.
L=S+0.5m=1.25m+0.5m=1.75m ANSWER 2 The long sides are 1.75 meter.
CHECK:
P=2(L+S)
6m=2(1.75m+1.25m)
6m=2(3m)
6m=6m